Please answer all the questions.

2) You have invested in a project that has the following payoff schedule:
Probability of
Payoff Occurrence
$40 .15
$50 .20
$60 .30
$70 .30
$80 .05
What is the expected value of the investment's payoff? (Round to the nearest $1.)
A) $60
B) $65 C) $59
D) $70

3) Spartan Sofas, Inc. is selling for $50.00 per share today. In one year, Spartan will be selling for $48.00 per share, and the dividend for the year will be $3.00. What is the cash return on Spartan stock?
A) $51.00 B) $1.00
C) $2.00
D) $3.00

4) You are considering investing in a firm that has the following possible outcomes:
Economic boom: probability of 25%; return of 25%
Economic growth: probability of 60%; return of 15%
Economic decline: probability of 15%; return of -5%
What is the standard deviation of returns on the investment?
A) 84.75%
B) 15.28%
C) 12.47% D) 9.21%

Roddy Richards invested $12,014.88 in Wolverine Meat Distributors (W.M.D.) five years ago. The investment had yearly arithmetic returns of -9.7%, -8.1%, 15%, 7.2%, and 15.4%.
6) What is the arithmetic average return of Roddy Richard's investment?
A) 2.42% B) 3.96%
C) 5.18%
D) 15.1%

7) What is the geometric average return of Roddy's Richard's investment? A) 3.38%
B) 4.63%
C) 6.96%
D) 8.78%

8) How much money did Roddy Richards receive when he sold his shares of W.M.D.?
A) $12,014.88
B) $12,398.42
C) $13,663.47 D) $14,184.73
Answer #1

1.
=40*0.15+50*0.20+60*0.30+70*0.30+80*0.05=59

2.
=48+3-50=1

3.
=sqrt(0.25*(25%-(0.25*25%+0.60*15%+0.15*(-5%)))^2+0.60*(15%-(0.25*25%+0.60*15%+0.15*(-5%)))^2+0.15*(-5%-(0.25*25%+0.60*15%+0.15*(-5%)))^2)
=9.206%

4.
=(-9.7%-8.1%+15%+7.2%+15.4%)/5=3.960%

5.
=((1-9.7%)*(1-8.1%)*(1+15%)*(1+7.2%)*(1+15.4%))^(1/5)-1=3.376%

6.
=12014.88*(1+3.376%)^5=14184.73